OpenStudy (anonymous):

A patients hospice cost stay 1/4 as much as his visit to the emergency room. His home nursing cost twice as much as his hospice stay. If his total healthcare bill was $140,000 how much did his home nursing cost. The answer the book came up with was 40,000 would someone break this down for me cause im lost the way the book explained it

OpenStudy (anonymous):

let x = cost of visit to emergency room then hospice cost x/4 home nursing = 2 * x/4 = x/2 form an equation: x + x/4 + x/2 = 140,000 solving for x 1 . 75 x = 140,000 x = 140,000 /1.75 = 80,000 dollars = emegency room cost home nursing = x/2 = 40,000 dollars

OpenStudy (anonymous):

(i'm assuming you know some algebra!)

OpenStudy (anonymous):

sorry - i should have asked you that ,first.

OpenStudy (anonymous):

I do but im not very strong at it but imma see if i can work this problem out right quick and ill jump online to ask u if i still need help or if i understand it

OpenStudy (anonymous):

hey jimmy rep i understand the beginnig part but i have copy and pasted the part I dont uderstand would u break these down just a tad bit more and explain sorry for beiing so difficult home nursing = 2 * x/4 = x/2 form an equation: x + x/4 + x/2 = 140,000 solving for x 1 . 75 x = 140,000 x = 140,000 /1.75 = 80,000 dollars = emegency room cost home nursing = x/2 = 40,000 dollars

OpenStudy (anonymous):

home nursing = 2x = 2x / 2 x -- ----- = -- 4 4 / 2 2 - divide top and bottom by 2 to simplify the fraction that help?

OpenStudy (anonymous):

ok give me one second here

OpenStudy (anonymous):

ok i got that part now on how you got the 2x but now how we get the answer 40,000